### v1

@@ 1 Introduction
In the 1960s, Tony Skyrme proposed a model of nucleons and atomic nuclei using topological solitons, now known as Skyrmions, within a low-energy effective theory of quantum chromodynamics [Skyrme:1961vq, Skyrme:1962vh]. Since then, Skyrmion-like topological solitons have been found not only in high-energy physics but also in a wide range of physical systems, including magnets [Mühlbauer:2009, Yu:2010ayi, Nagaosa:2013], liquid crystals [2014PhRvE..90a2505A], and optics [2018Sci...361..993T, 2019NatPh..15..650D, 2023LPRv...1700155M, 2024NaPho..18...15S]. In particular, two-dimensional analogues of Skyrmions in magnetic materials have been intensively investigated over the past fifteen years, following their experimental discovery [Mühlbauer:2009, Yu:2010ayi]. These 2D magnetic Skyrmions are also regarded as promising candidates for applications in nanotechnology.
Since the order parameter space (OPS) of (anti-) ferromagnets is $S^{2}=\mathbb{C}\mathrm{P}^{1}$, 2D Skyrmions in such a magnet are often referred to as $\mathbb{C}\mathrm{P}^{1}$ Skyrmion. As a natural generalization, 2D Skyrmions with the OPS being $\mathbb{C}\mathrm{P}^{N}=\mathrm{SU}(N+1)/\mathrm{U}(N)$, called $\mathbb{C}\mathrm{P}^{N}$ Skyrmion, have also been theoretically studied in several condensed matter systems [2008PhRvL.100d7203I, Garaud:2011zk, Garaud:2012pn, Kovrizhin:2012ff, Akagi:2021dpk, Zhang:2023, Benfenati:2022jvq, Amari:2022boe, 2025arXiv250917403L] [^1] The manifold $\mathbb{C}\mathrm{P}^{N}$ appears as the OPS of some physical systems. For instance, the OPS of $\frac{l}{2} spin$or Bose-Einstein Condensates is $\mathbb{C}\mathrm{P}^{l}$ [PhysRevB.81.184526], and that of $M$-layered Hall system is $\mathbb{C}\mathrm{P}^{2M-1}$ [PhysRevLett.82.3512]. . Note that, thanks to the homotopy group $\pi_{2}(\mathbb{C}\mathrm{P}^{N})=\mathbb{Z}$ for any positive integer $N$, $\mathbb{C}\mathrm{P}^{N}$ Skyrmions are characterized by an integer topological charge. Since $\mathbb{C}\mathrm{P}^{N}$ Skyrmions possess a large number of degrees of freedom, they reveal exotic pattern formation [Akagi:2021lva, Amari:2022boe] and intriguing dynamics [2025arXiv250917403L]. Such novel properties, absent in $\mathbb{C}\mathrm{P}^{1}$ Skyrmions, may pave new ways in the development of Skyrmion-based nanotechnology. In this article, we study the minimal case, i.e., $\mathbb{C}\mathrm{P}^{2}$ Skyrmions, for simplicity.
It is well known that topological solitons exhibit nontrivial spin–statistics properties determined by their topological charge [Wilczek:1983, Amari:2019tgs]. Like other cases, $\mathbb{C}\mathrm{P}^{2}$ Skyrmions support fermionic zero modes [Amari:2023gjq, Amari:2024rpm], of which number is equal to the topological invariant as shown by the Atiyah–Singer index theorem [Atiyah:1963zz]. Occupying these zero modes not only realizes the expected spin–statistics features but also induces a backreaction on the Skyrmion, thereby allowing one to investigate the interplay among fermion number, symmetry constraints, and soliton stability in a dynamical and self-consistent manner.
@@ 2 The Model
where $\mu^{2}>0$ and $U_{\infty}$ is the asymptotic value of the $U$ field. This potential is the square of the so-called old-baby potential. We employ this because the ordinary old-baby potential does not support $\mathbb{C}\mathrm{P}^{2}$ soliton solutions [PhysRevD.92.045007].
@@
It follows that the topological charge $Q$ depends on the boundary conditions of $F$ and $G$. Since we have $Z_{\infty}\propto\quantity(0,0,1)^{T}$, they must satisfy $F\quantity(\infty)=G\quantity(\infty)=\frac{k\pi}{2}\quantity(k\in\mathbb{Z})$. For the regularity of the static energy at the origin, $\sin F\quantity(0)$ must be 0, i.e., $F\quantity(0)=l\pi\quantity(l\in\mathbb{Z})$. Therefore, we employ $F\quantity(0)=G\quantity(0)=0,\ F\quantity(\infty)=G\quantity(\infty)=\frac{\pi}{2}$. Then, the topological charge $Q$ equals to $n_{2}$. In this article, we consider the case with $(n_{1},n_{2})=(1,3)$, giving $Q=3$, because it is pointed out in Ref. [PhysRevD.92.045007] that $Q=3$ solutions are the most simple nontrivial one in the Skyrme-type model (1) with the potential (2).
@@ 3 Numerical Results
One of the typical numerical results is presented in Fig.1, which shows the energy spectra for several choices of the coupling constant $m$. As one can see in Fig.1, there are three zero-crossing modes from negative to positive. This behavior is called spectral flow. These modes have corresponding quantum numbers $\kappa=\frac{5}{6},-\frac{1}{6},-\frac{7}{6}$ respectively. The number of zero-crossing modes equal to the topological charge $Q=3$ of the Skyrmion. This fact is a consequence of the Atiyah-Singer index theorem [Atiyah:1963zz]. Note that in order to input the fermion field into the fermion-Skyrmion interaction term of equation (7), it is necessary to choose the grand spin $\kappa$. Here we consider the case of $\kappa=\frac{5}{6}\quantity(\ell=-1)$.
@@
is localized around the Skyrmion center. Increasing the coupling constant $m$, the fermion density $\rho$ becomes strongly localized. In some previous studies, it is shown that this kind of localized fermions can deform the Skyrmion [Perapechka:2018yux, Dzhunushaliev:2024kti]. Such an effect is called backreaction effect. The bosonic field is shown in Fig.3. To see this, we compare the Skyrmion energy density $\mathcal{E}_{b}=-\mathcal{L}_{b}$ for weak and strong coupling constant $m$ in Fig. 4. The left panel shows the total bosonic energy density for $m=0.2,3.0$. The right panel plots the difference between the two densities $\Delta\equiv\mathcal{E}_{b}|_{m=3.0}-\mathcal{E}_{b}|_{m=0.2}.$ From this result, one sees that the backreaction makes the Skyrmion strongly localized. In other words, the Skyrmion-fermion coupling in our model (1) corresponds to an attractive force.
@@ 4 Summary and Outlook
When fermions are absent, the Skyrmion solution is stabilized by the balance between Skyrme term and potential term. Our results show that the fermion-Skyrmion interaction has a similar effect to the potential term, so it may be possible to replace the role of the potential term with the backreaction. In fact, $\mathbb{C}\mathrm{P}^{1}$ Skyrmions with localized fermions exist even in the absence of potential terms [Perapechka:2018yux]. We will analyze this possibility in the $\mathbb{C}\mathrm{P}^{2}$ case and reconsider the Derrick’s scaling argument [Derrick:1964] including isorotating fermions in the near future.

### v2

@@ 1 Introduction
In the 1960s, Tony Skyrme proposed a model of nucleons and atomic nuclei using topological solitons, now known as Skyrmions, within a low-energy effective theory of quantum chromodynamics [1, 2]. Since then, Skyrmion-like topological solitons have been found not only in high-energy physics but also in a wide range of physical systems, including magnets [3, 4, 5], liquid crystals [6], and optics [7, 8, 9, 10]. In particular, two-dimensional analogues of Skyrmions in magnetic materials have been intensively investigated over the past fifteen years, following their experimental discovery [3, 4]. These 2D magnetic Skyrmions are also regarded as promising candidates for applications in nanotechnology.
Since the order parameter space (OPS) of (anti-) ferromagnets is $S^{2}=\mathbb{C}\mathrm{P}^{1}$, 2D Skyrmions in such a magnet are often referred to as $\mathbb{C}\mathrm{P}^{1}$ Skyrmion. As a natural generalization, 2D Skyrmions with the OPS being $\mathbb{C}\mathrm{P}^{N}=\mathrm{SU}(N+1)/\mathrm{U}(N)$, called $\mathbb{C}\mathrm{P}^{N}$ Skyrmion, have also been theoretically studied in several condensed matter systems [11, 12, 13, 14, 15, 16, 17, 18, 19] [^1] The manifold $\mathbb{C}\mathrm{P}^{N}$ appears as the OPS of some physical systems. For instance, the OPS of $\frac{l}{2} spin$or Bose-Einstein Condensates is $\mathbb{C}\mathrm{P}^{l}$ [20], and that of $M$-layered Hall system is $\mathbb{C}\mathrm{P}^{2M-1}$ [21]. . Note that, thanks to the homotopy group $\pi_{2}(\mathbb{C}\mathrm{P}^{N})=\mathbb{Z}$ for any positive integer $N$, $\mathbb{C}\mathrm{P}^{N}$ Skyrmions are characterized by an integer topological charge. Since $\mathbb{C}\mathrm{P}^{N}$ Skyrmions possess a large number of degrees of freedom, they reveal exotic pattern formation [22, 18] and intriguing dynamics [19]. Such novel properties, absent in $\mathbb{C}\mathrm{P}^{1}$ Skyrmions, may pave new ways in the development of Skyrmion-based nanotechnology. In this article, we study the minimal case, i.e., $\mathbb{C}\mathrm{P}^{2}$ Skyrmions, for simplicity.
It is well known that topological solitons exhibit nontrivial spin–statistics properties determined by their topological charge [23, 24]. Like other cases, $\mathbb{C}\mathrm{P}^{2}$ Skyrmions support fermionic zero modes [25, 26], of which number is equal to the topological invariant as shown by the Atiyah–Singer index theorem [27]. Occupying these zero modes not only realizes the expected spin–statistics features but also induces a backreaction on the Skyrmion, thereby allowing one to investigate the interplay among fermion number, symmetry constraints, and soliton stability in a dynamical and self-consistent manner.
@@ 2 The Model
where $\mu^{2}>0$ and $U_{\infty}$ is the asymptotic value of the $U$ field. This potential is the square of the so-called old-baby potential. We employ this because the ordinary old-baby potential does not support $\mathbb{C}\mathrm{P}^{2}$ soliton solutions [28].
@@
It follows that the topological charge $Q$ depends on the boundary conditions of $F$ and $G$. Since we have $Z_{\infty}\propto\quantity(0,0,1)^{T}$, they must satisfy $F\quantity(\infty)=G\quantity(\infty)=\quantity(\frac{1}{2}+k)\pi\quantity(k\in\mathbb{Z})$. For the regularity of the static energy at the origin, $\sin F\quantity(0)$ must be 0, i.e., $F\quantity(0)=l\pi\quantity(l\in\mathbb{Z})$. Therefore, we employ $F\quantity(0)=G\quantity(0)=0,\ F\quantity(\infty)=G\quantity(\infty)=\frac{\pi}{2}$. Then, the topological charge $Q$ equals to $n_{2}$. In this article, we consider the case with $(n_{1},n_{2})=(1,3)$, giving $Q=3$, because it is pointed out in Ref. [28] that $Q=3$ solutions are the most simple nontrivial one in the Skyrme-type model (1) with the potential (2).
@@ 3 Numerical Results
One of the typical numerical results is presented in Fig.1, which shows the energy spectra for several choices of the coupling constant $m$. As one can see in Fig.1, there are three zero-crossing modes from negative to positive. This behavior is called spectral flow. These modes have corresponding quantum numbers $\kappa=\frac{5}{6},-\frac{1}{6},-\frac{7}{6}$ respectively. The number of zero-crossing modes equal to the topological charge $Q=3$ of the Skyrmion. This fact is a consequence of the Atiyah-Singer index theorem [27]. Note that in order to input the fermion field into the fermion-Skyrmion interaction term of equation (7), it is necessary to choose the grand spin $\kappa$. Here we consider the case of $\kappa=\frac{5}{6}\quantity(\ell=-1)$.
@@
is localized around the Skyrmion center. Increasing the coupling constant $m$, the fermion density $\rho$ becomes strongly localized. In some previous studies, it is shown that this kind of localized fermions can deform the Skyrmion [29, 30]. Such an effect is called backreaction effect. The bosonic field is shown in Fig.3. To see this, we compare the Skyrmion energy density $\mathcal{E}_{b}=-\mathcal{L}_{b}$ for weak and strong coupling constant $m$ in Fig. 4. The left panel shows the total bosonic energy density for $m=0.2,3.0$. The right panel plots the difference between the two densities $\Delta\equiv\mathcal{E}_{b}|_{m=3.0}-\mathcal{E}_{b}|_{m=0.2}.$ From this result, one sees that the backreaction makes the Skyrmion strongly localized. In other words, the Skyrmion-fermion coupling in our model (1) corresponds to an attractive force.
@@ 4 Summary and Outlook
When fermions are absent, the Skyrmion solution is stabilized by the balance between Skyrme term and potential term. Our results show that the fermion-Skyrmion interaction has a similar effect to the potential term, so it may be possible to replace the role of the potential term with the backreaction. In fact, $\mathbb{C}\mathrm{P}^{1}$ Skyrmions with localized fermions exist even in the absence of potential terms [29]. We will analyze this possibility in the $\mathbb{C}\mathrm{P}^{2}$ case and reconsider the Derrick’s scaling argument [31] including isorotating fermions in the near future.